OpenStudy (anonymous):

Justine enrolls in a music program. The cost of the program includes a one-time registration fee of $45 and a fee of $9 per lesson. If the average cost per music lesson is $12 and Justine has taken x classes, which equation represents this scenario?

OpenStudy (coconutjj):

is there choices?

OpenStudy (anonymous):

yes A.12=54/x B.12=45+9/x C.12=45+9x/x D.12=45x+9/x

OpenStudy (anonymous):

i think it is A.

OpenStudy (coconutjj):

It cannot be A, she could not have taken 4.5 classes right?

OpenStudy (anonymous):

ooh. true

OpenStudy (coconutjj):

question is a little comfusing. It gives you the average cost per music lesson, but doesn't say if it includes the registration fee. I'm tempted to say C but I have doubts too.

OpenStudy (anonymous):

the question is very vague, I'm confused but ill just put C. thank u

OpenStudy (anonymous):

that was right!!!